OpenStudy (anonymous):

Reduce [(96^2)(81^2)]/(216^3). Please help. Thank you.

OpenStudy (anonymous):

\[\Large \frac{(96^2)(81^2)}{(216^3)}\] Is this your qeustion?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Is 6 the answer? I'm not supposed to use a calculator, but I did.

OpenStudy (anonymous):

So how do I do this without a calculator?

terenzreignz (terenzreignz):

With cancellations. But factor everything out first.... (prime factors would be nice, btw)

OpenStudy (anonymous):

Yes first you should factor it out

OpenStudy (anonymous):

So this is basically a problem about division by hand?

OpenStudy (anonymous):

Basically, but sometimes you can have varaibles like a^-3/a^2 so this helps.

OpenStudy (anonymous):

OK. Should I try to divide everything by 3 first? Or is there a better prime number to try?

OpenStudy (anonymous):

You might want to try to do this first \[\Large at~the~topp~~is~(96^2)(81^2)=(81*96)^2\]

OpenStudy (anonymous):

Bottom is just 216^3

OpenStudy (anonymous):

\[\Large \frac{(96 \times 81)^2}{(216^3)}\]

OpenStudy (anonymous):

I can take it from here. Thank you all.

OpenStudy (anonymous):

Np, anytime you need help, just ask, thats what we're here for =)